Ch 14 Presentations &Homework eBook Calculator Measures of liquidity, Solvency and Profitability The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall Inc. common stock was $ 53 on December 31, 20Y2 Marshall Inc. Comparative Retained Earnings Statement For the Years Ended December 31, 20Y2 and 20Y1 20Y1 20Y2 Retained earnings, January 1 1,957,200 $1,664,600 414,000 341,000 Net income $2,371,200 $2,005,600 Total Dividends On preferred stock 7,000 7,000 41,400 41,400 On common stock 48,400 $48,400 Total dividends $2,322,800 1,957,200 Retained earnings, December 31
Ch 14 Presentations & Homework Calculator eBook Marshall Inc. Comparative Income Statement For the Years Ended December 31, 20Y2 and 20Y1 20Y1 20Y2 Sales $2,493,680 $ 2,297,590 789,800 858,480 Cost of goods sold $1,635,200 $ 1,507,790 Gross profit 670,390 Selling expenses $563,670 393,720 480,160 Administrative expenses Total operating expenses 1,043,830 1,064,110 $591,370 Income from operations $443,680 31,130 28,320 Other income 622,500 $ 472,000 84,000 152,000 Other expense (interest) $470,500 Income before income tax $ 388,000 47,000 56,500 Income tax expense $414,000 $341,000 Net income
Marshall Inc. Comparative Balance Sheet December 31, 20Y2 and 20Y1 Dec. 31, 20Y2 Dec. 31, 20Y1 Assets Current assets Cash 513,160 $411,900 Marketable securities 776,670 682,590 Accounts receivable (net) 459,900 430,700 262,800 Inventories 350,400 97,086 82,380 Prepaid expenses Total current assets $2,197,216 $1,870,370 Long-term investments 1,251,824 639,499 2,280,000 2,052,000 Property, plant, and equipment (net) $5,729,040 $4,561,869 Total assets Liabilities $ 646,240 $694,669 Current liabilities Long-term liabilities Mortgage note payable, 8 % $850,000 $ 0 1.050.000 1.050.000 Rnds nava hle, 8 %
Ch 14 Presentations &Homework eBook Calculator Mortgage note payable, 8 % 850,000 $ 0 1,050,000 1,050,000 Bonds payable, 8 % 1,900,000 1,050,000 Total long-term liabilities $2,546,240 $1,744,669 Total liabilities Stockholders' Equity 400,000 Preferred $ 0.70 stock, $40 par $400,000 Common stock, $ 10 par 460,000 460,000 2,322,800 1,957,200 Retained earnings $2,817,200 3,182,800 Total stockholders' equity $5,729,040 4,561,869 Total liabilities and stockholders' equity Required: Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year. 646,240x 1. Working capital 2. Current ratio Quick ratio 4. Accounts receivable turnover 5. Number of days' sales in receivables days 6. Inventory turnover
Ch 14 Presentations & Homework Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year 1. Working capital 2. Current ratio 3. Quick ratio 4. Accounts receivable turnover 5. Number of days sales in receivables days 6. Inventory turnover 7. Number of days' sales in inventory days 8. Ratio of fixed assets to long-term liabilities 9 Ratio of liabilities to stockholders' equity 10. Times interest earned 11. Asset turnover 12. Return on total assets 13. Return on stockholders' equity 14. Return on common stockholders' equity 15. Earnings per share on common stock 16. Price-earnings ratio 17. Dividends per share of common stock
Answer #1

1.Working Capital = Current Assets - Current Liabilities = 2197216 - 646240 = 1550976

2.Current Ratio = Current Assets / Current Liabilities = 2197216 / 646240 = 3.4

3.Quick Ratio = Quick Assets / Current Liabilities = (513160 + 776670 + 459900) / 646240 = 2.71

4.Accounts Receivable Turnover = Sales / Avg.Receivables = 2493680 / ( 459900 + 430700 ) / 2 = 2493680 / 445300 = 5.6

5.No. of days sales in receivables = 365 / Accounts Receivable Turnover = 365 / 5.6 = 65.18 Days

6.Inventory Turnover = Cost of goods sold / Avg.Inventory = 858480 / ( 350400 + 262800 ) / 2 = 858480 / 306600 = 2.8

7.No. of days sales in inventory= 365 / Inventory Turnover = 365 / 2.8 = 130.36 Days

8.Ratio of fixed assets to long term liabilities = 2280000 / 1900000 = 1.2

9.Ratio of Liabilities / Stockholders Equity = (646240 + 1900000 ) / 3182800 = 0.8

10.Times Interest Earned = EBIT / Interest Expense = (591370 + 31130 ) / 152000 = 4.095

11.Asset Turnover = Sales / Avg. Total Assets = 2493680 / (5729040 + 4561869 ) /2 = 2493680 / 5145455 = 0.485

12.Return on Total Assets = Net Income / Avg. Total Assets

                                      = 414000 / (5729040 + 4561869 ) /2 = 414000 / 5145455 = 8.05%

13.Return on Stockholders Equity = Net Income / Avg. Stockholders Equity

                                                  = 414000 / (3182800 + 2817200 ) /2 = 414000 / 3000000 = 13.8%

14.Return on Common Stockholders Equity = Net Income / Avg. Common Stockholders Equity

                                     = 414000 / (460000 + 460000 + 2322800 + 1957200 ) /2 = 414000 / 2600000 = 15.92%

15.Earnings per share on Common Stock = (Net Income - Preferred stock dividends) /No. of outstanding shares

                                     = (414000 - 7000 ) / 46000 = 407000 / 46000 = $8.85

16.Price Earnings Ratio = Market Price / EPS = $53 / $8.85 = 5.99

17.Dividend per share of Common Stock = Dividend on common stock / No. of outstanding shares

                                    = 41400 / 46000 = $0.90